OpenStudy (anonymous):

Bailey was mixing some instant pudding. The directions said to add 2 cups of milk to the contents of the package. The only measuring cup that she could find was a 2 3 cup. How many times should she fill this measuring cup? A) 2 B) 3 C) 4 D) 5

OpenStudy (anonymous):

3

OpenStudy (arabpride):

Alright for this one you have to set up your OWN equation to solve.. so.. 2/3x = 3 and we solve....

OpenStudy (arabpride):

<set it up as..> 2/3x = 3/1

OpenStudy (anonymous):

9/2

OpenStudy (dmndlife24):

I think she meant 2/3x=2

OpenStudy (anonymous):

then its 3???

OpenStudy (arabpride):

Actually - allow me to correct myself~ 3 ÷ 2/3

OpenStudy (arabpride):

Leave the question and come back in to see clearly plz~

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The answer is C. 4

OpenStudy (arabpride):

okay since we're dividing... we get the equation 3/1 * 3/2 = 9/2 ~like you said earlier~

OpenStudy (arabpride):

Now change it from an improper fraction~

OpenStudy (dmndlife24):

 This is a very simple way to look at it